Can you sketch the graph of this line: 8x+5y=-20

Answer by ewatrrr(24785)


Hi
sketch the graph of this line:
8x+5y=-20 x = 0 then y =-4 AND when y = 0 then x = -20/8 = -5/2
Pt(0,-4) and Pt(-5/2,0) on the line. Plot and connect.